Sketch the region enclosed by the curves and find its area.



y=e^x, y=e^8x, x=0, x=ln2

1
Expert's answer
2022-01-26T18:09:31-0500
"e^x=e^{8x}=>x=0"
"A=\displaystyle\int_{0}^{\ln 2}(e^{8x}-e^x)dx=[\dfrac{e^{8x}}{8}-e^x]\begin{matrix}
 \ln 2 \\
 0
\end{matrix}"

"=\dfrac{256}{8}-2-(\dfrac{1}{8}-1)"

"=\dfrac{247}{8}({units}^2)=30.875{units}^2"